OpenStudy (anonymous):

What is the value of M? 4/9 = m/54 ?

OpenStudy (a1234):

Cross multiply and solve. 4 * 54 = 216 216/9 = 24 m = 24

OpenStudy (whpalmer4):

\[\frac{4}{9} = \frac{m}{54}\]Cross-multiply to solve: \[4*54 = m*9\]\[\frac{4*54}{9} = m =24 \]

OpenStudy (anonymous):

I did cross multiply and i got 24 but it said it was wrong?

OpenStudy (whpalmer4):

Did you copy the problem correctly? \[\frac{4}{9}*\frac{6}{6} = \frac{24}{54}\]

OpenStudy (jdoe0001):

I did cross multiply and i got 24 \(\large \checkmark\)

OpenStudy (anonymous):

I know i got 24 too the first time i did it i dont understand how i got it wrong ? ..

OpenStudy (whpalmer4):

Whatever your system says, you can take comfort in knowing that you correctly found the value of \(m\) in the problem you showed us. If you typed the answer wrong, or the system had the wrong answer programmed, or it was supposed to be a slightly different problem, well, whatever. You did this one right as far as we're concerned :-)

OpenStudy (a1234):

Unless there is a problem with what you typed, 24 is correct.

OpenStudy (anonymous):

Oh okay thanks :)